I need help please, I cannot figure this out. thank you.
At a large college, the probability that students are from Ohio is 4%. A random sample of 20 students from the college is selected. Let X denote the number of students from Ohio among them.
Using the Normal Approximation to the Binomial Distribution...
The mean of X is:
The standard deviation of X is:
The probability that X is at least 1 is:

Answer by stanbon(75887)

At a large college, the probability that students are from Ohio is 4%.
A random sample of 20 students from the college is selected.
Let X denote the number of students from Ohio among them.
----------------------------------------------------------------
Using the Normal Approximation to the Binomial Distribution...
The mean of X is: np = 20*0.04 = 0.8
--------------------------------------------
The standard deviation of X is: sqrt(npq) = sqrt(0.8*0.96) = 0.8764
--------------------------------------------
The probability that X is at least 1 is:
P(X at least 1) = 1 - P(X = 0) = 1-(0.96)^20 = 0.558
=====================================
Cheers,
Stan H.
===================